# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

### v4

@@ A.4 Text Classification
Table 2 shows commonly used datasets (e.g. HoC and OHSUMED) for text classification. Since this is a text-level task, the [CLS] vector is augmented when using task-specific fine-tuned (Transformer-based) models. Another approach for distilling the overall information is to use the weighted sum of final attention layer outputs. Recent work has shown that adding a custom attention layer after the original model (BERT) improves the model performance [260, 261]. McCreery et al. [262] double-fine-tuned the BERT model, in the sense that they first fine-tuned the model on a general dataset and then fine-tuned it on a medical-specific dataset. Another approach is to combine graph-based models with general LLMs, as in ChatGraph [258] that used ChatGPT to extract text information and apply it to a graph-based model that outperforms GPT models. CohortGPT [263] used Chain-of-Thought prompting and knowledge graph to outperform few-shot ChatGPT and GPT-4.
@@
Same as above, the F1-score is used for evaluating the model performance. Table 6 compares the performance of PubMedBERT [30], BioGPT [94], GPT-4 [35, 263], ChatGPT [258, 263], and ChatGraph [258] on four benchmark datasets. It shows that GPT models perform better on the i2b2 dataset than on the OHSUMED dataset, possibly because its few-shot setting is not yet sufficiently good. Overall, few-shot GPT models outperform zero-shot ones, and the F1 scores tend to increase with more parameters.
@@ A.5 Natural Language Inference
The commonly used datasets, such as MedNLI [165] and BioNLI [166], are shown in Table 2 . Due to the scarcity of biomedical datasets, researchers also use general datasets (e.g. SNLI [264] and MultiNLI [265]). In task-specific fine-tuned models, similar to RE, [CLS] vector is often added to the text to show the overall information. The difference is that here we have two (instead of one) sentences, and so a [SEP] vector is also applied to indicate the separation of premise and hypothesis, yielding an overall input that looks like ‘[CLS], premise, [SEP], hypothesis’. For BERT-based models, a classification head is applied to the final [CLS] vector to predict the label. Kanakarajan et al. [90] first pre-trained BioBERT [29] on MIMIC-III [39] and then fine-tuned the model on MedNLI [165]. Cengiz et al. [266] used the so-called two-stage sequential transfer learning method. They trained the BioBERT model on SNLI [264] and MultiNLI [265] first then fine-tuned it on the MedNLI dataset [165]. The majority vote method is used to combine the prediction output of different trained models. Gu et al. [248] combined GPT-3.5 and PubMedBERT using knowledge distillation. They first fed GPT-3.5 original texts and asked it to distill the data, and the distilled data was subsequently used to further train the PubMedBERT model. Wang et al. [233] defined the self-questioning prompting (as discussed in Sec A.1) and utilize it in the MedNLI task.
@@
The F1-score is used in the same manner as above. Table 7 compares the performance of task-specific fine-tuned models (e.g. BioBERT, ClinicalBERT, BlueBERT, and SciBER) on the benchmark MedNLI dataset [248], together with the performances of GPT-3.5 and GPT-3.5-Distillation. It shows that BioBERT [29] has the best performance, and the current general LLMs still need further improvement on natural language inference.
@@ A.6 Semantic Textual Similarity
As shown in Table 2, the ‘2019 n2c2/OHNLP’ [168], BIOSSES [169], and MedSTS [167] are widely-used benchmark datasets for STS. Since STS is a text-level task, [CLS] vector is naturally added to extract the overall information. Mutinda et al. [227] added a fully connected layer after the [CLS] vector as the architecture. Yang et al. [80] combined the representation of different models and added a fully connected layer. Wang et al. [267] dropped the concept of [CLS] and used the Hierarchical Convolution (HConv) layer as the added last layer. Wang et al. [233] defined the self-questioning prompting (SQP) and utilize it on the BIOSSES dataset. Xiong et al. [268] used the concatenation of character level, sentence level, and entity level representation (CSE-concate) to extract the information that is fed to the further added MLP layer.
For further training on a pre-trained model, there is a common understanding that the new dataset should have a similar distribution to that of the dataset for pre-training to improve the model performance. The iterative training is introduced for this purpose [269, 228], where the phrase "iterative" refers to the following two iterative steps: a) It first freezes the model and computes the outputs from the dataset, then it chooses a subset of the dataset so the outputs of the subset have a similar distribution to that of the pre-training dataset. b) The model is further trained using the obtained subset obtained in Step a).
@@
The Pearson correlation coefficient $r_{xy}$ is used for evaluation purposes. For $n$ pairs of data $\{x_{i},y_{i}\}_{i=1}^{n}$, we define $r_{xy}=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}}\,.$ Here, $\{x_{i}\}_{i=1}^{n}$ represents the actual similarity and $\{y_{i}\}_{i=1}^{n}$ represents the predicted similarity. Table 8 compares the performance of BERT [227, 257], ClinicalBERT [227, 257], XLNet [80], RoBERTa [80, 257], HConv-BERT [267], BARD [233], GPT-3.5 [233], GPT-4 [233], BERT (CSE-concate) [268], and ClinicalBERT (iterative training) [228] on BIOSSES and ‘2019 n2c2/OHNLP’ datasets.
@@
Task-specific FT Models BERT [70] 81.40 69.23
ClinicalBERT [31] 91.23 83.20
XLNet [80] - 84.70
RoBERTa [80, 81] 81.25 87.78
HConv-BERT [267] - 79.40
BERT (CSE-concate) [268] - 86.80
ClinicalBERT (iterative training) [31] - 87.00
General LLMs BARD (zero-shot) [87] 57.60 -
BARD (5-shot) [87] 60.10 -
GPT-3.5 (zero-shot) [6] 87.30 -
GPT-3.5 (5-shot) [6] 89.20 -
GPT-4 (zero-shot) [7] 88.90 -
GPT-4 (5-shot) [7] 91.60 -
@@
Task-specific FT Models OpenAI cpt-text-S [270, 93] 67.9 33.2 -
OpenAI cpt-text-M [270, 93] 58.5 36.7 -
OpenAI cpt-text-L [270, 93] 56.2 38.0 -
OpenAI cpt-text-XL [270, 93] 64.9 40.7 -
BioCPT [29, 90] 70.9 35.5 55.3
General LLMs GTR-Base [271] 53.9 30.8 27.1
GTR-Large [271] 55.7 32.9 32.0
GTR-XL [271] 58.4 34.3 31.7
GTR-XXL [271] 50.1 34.2 32.4
ChatGPT [239] 76.7 35.6 -
ChatGPT [239] + GPT-4 [7] 85.5 38.5 -
@@ A.7 Information Retrieval
Table 2 shows commonly used datasets, such as TREC-COVID [170], NFCorpus [171], and BioASQ [172] in the biomedical area. In addition, a general dataset, BEIR (Benchmarking IR) [172], is often included in the training step of biomedical IR. Jin et al.
[93] developed a BERT-based model BioCPT that encodes the query and articles for the ranking task. They split the method into training, inference, and evaluation steps. In the training step, they introduce query-to-document loss and document-to-query loss to train the encoders for the query and articles. In the inference step, they concatenate the encoding of the query and its best-fit article $d_{1}^{q}$ together with $k-1$ non-relevant articles $(d_{2}^{q},d_{3}^{q},\cdots,d_{k}^{q})$ found by maximum inner product search (MIPS) to further train the model to rank $d_{1}^{q}$ at the top. In the evaluation step, for each input query $q$, the model evaluates over the whole dataset $\{d_{i}\}_{i=1}^{n}$ to find the best $k$ relevant articles by MIPS and used the model from the inference step to rank those $k$ articles.
Sun et al. [272] applied zero-shot ChatGPT to rank the most relevant documents without abnormal prompting. They also further used GPT-4 to re-rank the top 30 documents retrieved by ChatGPT. Abonizio et al. [273] introduced two LLM-based data augmentation methods, namely InPars and Promptagator, for IR. For InPars, they used GPT-3 and GPT-J to generate a new query for a randomly selected document. They used few-shot prompting which provides the model with good query examples or bad query examples. For Promptagator, the major difference is that a more dataset-specific prompting is applied. Ateia and Kruschwitz [274] proposed a query expansion technique that expands the current query into a more comprehensive query, which consistently improves the performance of any successive tasks. It is done purely by the GPT model with regular instructional prompting. Similarly, Wang et al. [275] used ChatGPT to generate more refined Boolean queries for systematic reviews. They showed that ChatGPT is able to generate or refine queries with higher precision.
@@
We compare the performance of Google’s Generalizable T5-based dense Retrievers (GTR) [276, 93], OpenAI’s cpt-text [270, 93], BioCPT [93], ChatGPT [272], and ChatGPT+GPT-4 [272] on the datasets we introduced before.
@@ B.1 Text Summarization
Table 2 shows the commonly used biomedical datasets PubMed [38] and MentSum [174]. Some general datasets, such as GigaWord [277], CL-SciSumm [278], and S2ORC [110], are also commonly used for text summarization.
For extractive summarization, the key task is to define some scoring system that scores all sentences and hence finds the most important ones. Moradi et al. [279] used a clustering-based method to summarize medical text. They vectorize the tokens of the text by BERT and cluster the sentence vector into $k$ clusters, and then they define an informativeness score that chooses one sentence from each cluster to form a summary. Moradi et al. [280] used the graph-based model for summarization, where sentences are treated as nodes and relations as edges. The relations are measured by the cosine similarity of vectors representing the sentences. They then used different graph ranking algorithms to choose important sentences as a summary.
Du et al. [281] used purely the Transformer-based model as the scoring system. They i) tokenized the whole text with [CLS] and [SEP] augmented, ii) further augmented the corresponding sentence and token positions to each token, and iii) fed the whole vector into the model. A sigmoid layer is added to the model so the output is between zero and one, and the output is considered the score for each sentence. Since the Transformer automatically extracts relations of one sentence to others, there is no need for manually designing a score. Chen et al. [282] also relied only on the Transformer to score the sentences. Different from Du et al. [281], they proposed the AlphaBERT that takes the character-level tokens as input,and the training is split into pre-training and fine-tuning.
McInerney et al. [283] combined the summarization model with a query to generate specific summaries. Pang et al. [284] proposed a principled inference framework with top-down and bottom-up inference techniques to improve summarization models.
There are also studies on summarizing more than one piece of text at a time. They mainly relied on graph-based or Transformer-based models to extract relations between different texts [285]. Zhang et al. [286] applied one-shot GPT-3.5, using dialogues and summaries from the same category as prompts to generate abstractive summarization. More studies are needed to qualitatively analyze the performance of GPT models on biomedical text summarization [287].
@@
The standard Recall-Oriented Understudy for Gisting Evaluation (ROUGE) data[288], including ROUGE-1 and ROUGE-2, is used for evaluation. Table 10 compares the performance of BioBERT variations using different training sets [280], rankings (PageRank and PPF) [280], and added last layers [281] on PubMed. It shows that models fine-tuned on Pubmed or PMC data achieve improved performance. Differences in the ranking strategy do not correspond to appreciable differences in the model performance. Overall, current BERT models are far below the expectation of true text summarization, measured by the ROUGE-2 scores.
@@
Task-specific FT Models BioBERT (LSTM) [281] 35.82 & 17.15
BioBERTSum [281] 37.45 & 17.59
BioBERT (Pubmed + PPF) [280] 74.21 & 32.88
BioBERT (PMC + PageRank) [280] 75.82 & 34.01
BioBERT (Pubmed + PageRank) [280] 76.09 & 34.38
BioBERT (Pubmed + PMC + PageRank) [280] 76.34 & 34.67
Table 11: The performance (BLEU score) of the text simplification task on the MultiCochrane (English) dataset, shown in percentage (%). Higher is better for BLEU.
@@
Model MultiCochrane
GPT-3 (zero-shot) [6] 2.38
Flan-T5 (zero-shot) [76] 8.12
mT5 [76] 8.82
Flan-T5 (fine-tuned) [76] 8.70
@@
BERT [70] 62.40
RoBERTa [80, 81] 53.28
XLNet [80] 46.20
GPT-2 [82] 49.00
BERT [70] + RoBERTa [80, 81] + XLNet [80] + GPT-2 [82] 64.52

### v5

@@ A.4 Text Classification
Table 4 shows commonly used datasets (e.g. HoC and OHSUMED) for text classification. Since this is a text-level task, the [CLS] vector is augmented when using task-specific fine-tuned (Transformer-based) models. Another approach for distilling the overall information is to use the weighted sum of final attention layer outputs. Recent work has shown that adding a custom attention layer after the original model (BERT) improves the model performance [tang2019progress, wood2020automated]. McCreery et al. [mccreery2019domain] double-fine-tuned the BERT model, in the sense that they first fine-tuned the model on a general dataset and then fine-tuned it on a medical-specific dataset. Another approach is to combine graph-based models with general LLMs, as in ChatGraph [shi2023chatgraph] that used ChatGPT to extract text information and apply it to a graph-based model that outperforms GPT models. CohortGPT [guan2023cohortgpt] used Chain-of-Thought prompting and knowledge graph to outperform few-shot ChatGPT and GPT-4.
@@
Same as above, the F1-score is used for evaluating the model performance. Table 8 compares the performance of PubMedBERT [gu2021domain], BioGPT [luo2022biogpt], GPT-4 [liu2023deid, guan2023cohortgpt], ChatGPT [shi2023chatgraph, guan2023cohortgpt], and ChatGraph [shi2023chatgraph] on four benchmark datasets. It shows that GPT models perform better on the i2b2 dataset than on the OHSUMED dataset, possibly because its few-shot setting is not yet sufficiently good. Overall, few-shot GPT models outperform zero-shot ones, and the F1 scores tend to increase with more parameters.
@@ A.5 Natural Language Inference
The commonly used datasets, such as MedNLI [romanov2018lessons] and BioNLI [bastan2022bionli], are shown in Table 4 . Due to the scarcity of biomedical datasets, researchers also use general datasets (e.g. SNLI [bowman2015large] and MultiNLI [williams2017broad]). In task-specific fine-tuned models, similar to RE, [CLS] vector is often added to the text to show the overall information. The difference is that here we have two (instead of one) sentences, and so a [SEP] vector is also applied to indicate the separation of premise and hypothesis, yielding an overall input that looks like ‘[CLS], premise, [SEP], hypothesis’. For BERT-based models, a classification head is applied to the final [CLS] vector to predict the label. Kanakarajan et al. [raj2019saama] first pre-trained BioBERT [lee2020biobert] on MIMIC-III [johnson2016mimic] and then fine-tuned the model on MedNLI [romanov2018lessons]. Cengiz et al. [cengiz2019ku_ai] used the so-called two-stage sequential transfer learning method. They trained the BioBERT model on SNLI [bowman2015large] and MultiNLI [williams2017broad] first then fine-tuned it on the MedNLI dataset [romanov2018lessons]. The majority vote method is used to combine the prediction output of different trained models. Gu et al. [gu2023distilling] combined GPT-3.5 and PubMedBERT using knowledge distillation. They first fed GPT-3.5 original texts and asked it to distill the data, and the distilled data was subsequently used to further train the PubMedBERT model. Wang et al. [wang2023large] defined the self-questioning prompting (as discussed in Sec A.1) and utilize it in the MedNLI task.
@@
The F1-score is used in the same manner as above. Table 9 compares the performance of task-specific fine-tuned models (e.g. BioBERT, ClinicalBERT, BlueBERT, and SciBER) on the benchmark MedNLI dataset [gu2023distilling], together with the performances of GPT-3.5 and GPT-3.5-Distillation. It shows that BioBERT [lee2020biobert] has the best performance, and the current general LLMs still need further improvement on natural language inference.
@@ A.6 Semantic Textual Similarity
As shown in Table 4, the ‘2019 n2c2/OHNLP’ [wang20202019], BIOSSES [souganciouglu2017biosses], and MedSTS [wang2020medsts] are widely-used benchmark datasets for STS. Since STS is a text-level task, [CLS] vector is naturally added to extract the overall information. Mutinda et al. [mutinda2020detecting] added a fully connected layer after the [CLS] vector as the architecture. Yang et al. [yang2020measurement] combined the representation of different models and added a fully connected layer. Wang et al. [wang2020evaluating] dropped the concept of [CLS] and used the Hierarchical Convolution (HConv) layer as the added last layer. Wang et al. [wang2023large] defined the self-questioning prompting (SQP) and utilize it on the BIOSSES dataset. Xiong et al. [xiong2020using] used the concatenation of character level, sentence level, and entity level representation (CSE-concate) to extract the information that is fed to the further added MLP layer.
For further training on a pre-trained model, there is a common understanding that the new dataset should have a similar distribution to that of the dataset for pre-training to improve the model performance. The iterative training is introduced for this purpose [wang2020learning, mahajan2020identification], where the phrase "iterative" refers to the following two iterative steps: a) It first freezes the model and computes the outputs from the dataset, then it chooses a subset of the dataset so the outputs of the subset have a similar distribution to that of the pre-training dataset. b) The model is further trained using the obtained subset obtained in Step a).
@@
The Pearson correlation coefficient $r_{xy}$ is used for evaluation purposes. For $n$ pairs of data $\{x_{i},y_{i}\}_{i=1}^{n}$, we define $r_{xy}=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}}\,.$ Here, $\{x_{i}\}_{i=1}^{n}$ represents the actual similarity and $\{y_{i}\}_{i=1}^{n}$ represents the predicted similarity. Table 10 compares the performance of BERT [mutinda2020detecting, wang2023pre], ClinicalBERT [mutinda2020detecting, wang2023pre], XLNet [yang2020measurement], RoBERTa [yang2020measurement, wang2023pre], HConv-BERT [wang2020evaluating], BARD [wang2023large], GPT-3.5 [wang2023large], GPT-4 [wang2023large], BERT (CSE-concate) [xiong2020using], and ClinicalBERT (iterative training) [mahajan2020identification] on BIOSSES and ‘2019 n2c2/OHNLP’ datasets.
@@
Task-specific FT Models BERT [devlin2018bert] 81.40 69.23
ClinicalBERT [alsentzer2019publicly] 91.23 83.20
XLNet [yang2020measurement] - 84.70
RoBERTa [yang2020measurement, van2020automets] 81.25 87.78
HConv-BERT [wang2020evaluating] - 79.40
BERT (CSE-concate) [xiong2020using] - 86.80
ClinicalBERT (iterative training) [alsentzer2019publicly] - 87.00
PubMedBERT [gu2021domain] 92.30 -
BioLinkBERT [yasunaga2022linkbert] 93.63 -
General LLMs ChatGPT [openai2023chatgpt] 43.75 -
BARD (zero-shot) [manyika2023overview] 57.60 -
BARD (few-shot) [manyika2023overview] 60.10 -
GPT-3.5 (zero-shot) [GPT-3] 87.86 -
GPT-3.5 (few-shot) [GPT-3] 91.94 -
GPT-4 (zero-shot) [gpt-4] 88.90 -
GPT-4 (few-shot) [gpt-4] 91.60 -
@@
Task-specific FT Models OpenAI cpt-text-S [neelakantan2022text, jin2023biocpt] 67.9 33.2 -
OpenAI cpt-text-M [neelakantan2022text, jin2023biocpt] 58.5 36.7 -
OpenAI cpt-text-L [neelakantan2022text, jin2023biocpt] 56.2 38.0 -
OpenAI cpt-text-XL [neelakantan2022text, jin2023biocpt] 64.9 40.7 -
BioCPT [lee2020biobert, raj2019saama] 70.9 35.5 55.3
General LLMs GTR-Base [google2023GTR] 53.9 30.8 27.1
GTR-Large [google2023GTR] 55.7 32.9 32.0
GTR-XL [google2023GTR] 58.4 34.3 31.7
GTR-XXL [google2023GTR] 50.1 34.2 32.4
ChatGPT [openai2023chatgpt] 76.7 35.6 -
ChatGPT [openai2023chatgpt] + GPT-4 [gpt-4] 85.5 38.5 -
@@ A.7 Information Retrieval
Table 4 shows commonly used datasets, such as TREC-COVID [voorhees2021trec], NFCorpus [boteva2016full], and BioASQ [thakur2021beir] in the biomedical area. In addition, a general dataset, BEIR (Benchmarking IR) [thakur2021beir], is often included in the training step of biomedical IR. Jin et al.
[jin2023biocpt] developed a BERT-based model BioCPT that encodes the query and articles for the ranking task. They split the method into training, inference, and evaluation steps. In the training step, they introduce query-to-document loss and document-to-query loss to train the encoders for the query and articles. In the inference step, they concatenate the encoding of the query and its best-fit article $d_{1}^{q}$ together with $k-1$ non-relevant articles $(d_{2}^{q},d_{3}^{q},\cdots,d_{k}^{q})$ found by maximum inner product search (MIPS) to further train the model to rank $d_{1}^{q}$ at the top. In the evaluation step, for each input query $q$, the model evaluates over the whole dataset $\{d_{i}\}_{i=1}^{n}$ to find the best $k$ relevant articles by MIPS and used the model from the inference step to rank those $k$ articles.
Sun et al. [sun2023chatgpt] applied zero-shot ChatGPT to rank the most relevant documents without abnormal prompting. They also further used GPT-4 to re-rank the top 30 documents retrieved by ChatGPT. Abonizio et al. [abonizio2023inpars] introduced two LLM-based data augmentation methods, namely InPars and Promptagator, for IR. For InPars, they used GPT-3 and GPT-J to generate a new query for a randomly selected document. They used few-shot prompting which provides the model with good query examples or bad query examples. For Promptagator, the major difference is that a more dataset-specific prompting is applied. Ateia and Kruschwitz [ateia2023chatgpt] proposed a query expansion technique that expands the current query into a more comprehensive query, which consistently improves the performance of any successive tasks. It is done purely by the GPT model with regular instructional prompting. Similarly, Wang et al. [wang2023can] used ChatGPT to generate more refined Boolean queries for systematic reviews. They showed that ChatGPT is able to generate or refine queries with higher precision.
@@
We compare the performance of Google’s Generalizable T5-based dense Retrievers (GTR) [ni2021large, jin2023biocpt], OpenAI’s cpt-text [neelakantan2022text, jin2023biocpt], BioCPT [jin2023biocpt], ChatGPT [sun2023chatgpt], and ChatGPT+GPT-4 [sun2023chatgpt] on the datasets we introduced before.
@@ B.1 Text Summarization
Table 4 shows the commonly used biomedical datasets PubMed [PubMed] and MentSum [sotudeh2022mentsum]. Some general datasets, such as GigaWord [napoles2012annotated], CL-SciSumm [yasunaga2019scisummnet], and S2ORC [lo2019s2orc], are also commonly used for text summarization.
For extractive summarization, the key task is to define some scoring system that scores all sentences and hence finds the most important ones. Moradi et al. [moradi2020deep] used a clustering-based method to summarize medical text. They vectorize the tokens of the text by BERT and cluster the sentence vector into $k$ clusters, and then they define an informativeness score that chooses one sentence from each cluster to form a summary. Moradi et al. [moradi2020summarization] used the graph-based model for summarization, where sentences are treated as nodes and relations as edges. The relations are measured by the cosine similarity of vectors representing the sentences. They then used different graph ranking algorithms to choose important sentences as a summary.
Du et al. [du2020biomedical] used purely the Transformer-based model as the scoring system. They i) tokenized the whole text with [CLS] and [SEP] augmented, ii) further augmented the corresponding sentence and token positions to each token, and iii) fed the whole vector into the model. A sigmoid layer is added to the model so the output is between zero and one, and the output is considered the score for each sentence. Since the Transformer automatically extracts relations of one sentence to others, there is no need for manually designing a score. Chen et al. [chen2020modified] also relied only on the Transformer to score the sentences. Different from Du et al. [du2020biomedical], they proposed the AlphaBERT that takes the character-level tokens as input,and the training is split into pre-training and fine-tuning.
McInerney et al. [mcinerney2020query] combined the summarization model with a query to generate specific summaries. Pang et al. [pang2022long] proposed a principled inference framework with top-down and bottom-up inference techniques to improve summarization models.
There are also studies on summarizing more than one piece of text at a time. They mainly relied on graph-based or Transformer-based models to extract relations between different texts [chen2021capturing]. Zhang et al. [zhang2023ds4dh] applied one-shot GPT-3.5, using dialogues and summaries from the same category as prompts to generate abstractive summarization. More studies are needed to qualitatively analyze the performance of GPT models on biomedical text summarization [shaib2023summarizing].
@@
The standard Recall-Oriented Understudy for Gisting Evaluation (ROUGE) data[lin2004rouge], including ROUGE-1 and ROUGE-2, is used for evaluation. Table 12 compares the performance of BioBERT variations using different training sets [moradi2020summarization], rankings (PageRank and PPF) [moradi2020summarization], and added last layers [du2020biomedical] on PubMed. It shows that models fine-tuned on Pubmed or PMC data achieve improved performance. Differences in the ranking strategy do not correspond to appreciable differences in the model performance. Overall, current BERT models are far below the expectation of true text summarization, measured by the ROUGE-2 scores.
@@
Task-specific FT Models BioBERT (LSTM) [du2020biomedical] 35.82 & 17.15
BioBERTSum [du2020biomedical] 37.45 & 17.59
BioBERT (Pubmed + PPF) [moradi2020summarization] 74.21 & 32.88
BioBERT (PMC + PageRank) [moradi2020summarization] 75.82 & 34.01
BioBERT (Pubmed + PageRank) [moradi2020summarization] 76.09 & 34.38
BioBERT (Pubmed + PMC + PageRank) [moradi2020summarization] 76.34 & 34.67
GPT-3.5 (zero-shot) [GPT-3] 6.08 & -
GPT-3.5 (few-shot) [GPT-3] 23.20 & -
GPT-4 (zero-shot) [gpt-4] 39.97 & -
GPT-4 (few-shot) [gpt-4] 40.54 & -
Table 13: The performance (BLEU score) of the text simplification task on the MultiCochrane (English) dataset, shown in percentage (%). Higher is better.
Model MultiCochrane
GPT-3 (zero-shot) [GPT-3] 2.38
Flan-T5 (zero-shot) [joseph2023multilingual] 8.12
mT5 [joseph2023multilingual] 8.82
Flan-T5 (fine-tuned) [joseph2023multilingual] 8.70
GPT-3.5 (zero-shot) [GPT-3] 13.05
GPT-3.5 (few-shot) [GPT-3] 13.18
GPT-4 (zero-shot) [gpt-4] 12.00
GPT-4 (few-shot) [gpt-4] 13.12
@@
BERT [devlin2018bert] 62.40
RoBERTa [yang2020measurement, van2020automets] 53.28
XLNet [yang2020measurement] 46.20
GPT-2 [gpt-2] 49.00
BERT [devlin2018bert] + RoBERTa [yang2020measurement, van2020automets] + XLNet [yang2020measurement] + GPT-2 [gpt-2] 64.52